[Hallc_running] [New Logentry] Status Update

pooser at jlab.org pooser at jlab.org
Sat Feb 24 01:35:02 EST 2018

Logentry Text:
There are still 2 owl shifts that require shift leaders, Monday and Tuesday (2/26 & 2/27).  Please sign up for these shifts if you can.

Day shift started out with the SHMS at -2.5 Gev, 39 deg, and HMS at -3.3 GeV, 21 deg with an LD2 target.  Shortly after 9 am, the HMS momentum was decreased to -1.5 Gev and swung out to 59 deg.  Production data was taken until ~1 pm at which time Q2 tripped (https://logbooks.jlab.org/entry/3537163).  This couldn’t have occurred at a more convenient time since the accelerator was taken down in order to replace the cooling fan on the C100 cavity 1L22.

At this time a controlled access was taken by the Steve L. et. al. to address the infamous Q2 power supply.  The magnet work was completed around 4pm (https://logbooks.jlab.org/entry/3537306) and 50 uA on LH2 was restored.  Issues with the HMS pre-shower and shower sum signals were resolved and the troublesome pedestals remained calm (https://logbooks.jlab.org/entry/3537466).

The remainder of swing shift was comprised of ratty beam (https://logbooks.jlab.org/entry/3537467) and production data on LH2, LD2, and 10 cm dummy targets according to the F2/XEM run plan were taken.  MCC increased our ramp rate in attempt to reduce the amount of downtime as a result of frequent trips.

Moving forward we will complete collecting the 10k e-’s with the 10 cm dummy target and then we will proceed with LH2 data taking.  We plan to stay there through the night so as to minimize configuration changes.
Work currently planned during Monday’s (2/26) beam studies starting at 9 am:
Joe Beaufait: diagnose (and potentially fix) HMS door/shutter breaker (https://logbooks.jlab.org/entry/3535292)
Andrew: install five Y strainers on the SHMS power supply LCW lines.  The power supplies will have to be ramped down and turned off

Special notices in effect for shift crews:
Trouble shooting HMS Q2 power supply: https://logbooks.jlab.org/entry/3537306
Don't ramp HMS Q2 higher than 610A: https://logbooks.jlab.org/entry/3535714
TO Instructions for LH2 <--> LD2 switches : https://logbooks.jlab.org/entry/3534487
Maximum Boron targets is 40 muA: https://logbooks.jlab.org/entry/3535297
Any missing reference time messages should not be ignored: https://logbooks.jlab.org/entry/3535289


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3537564
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180224/9f132494/attachment.html>

More information about the Hallc_running mailing list